A 2.50-L volume of hydrogen measured at –196 °C is warmed to 100 °C. Calculate the volume of the
gas at the higher temperature, assuming no change in pressure.
1
Expert's answer
2019-04-17T04:16:43-0400

According to the ideal gas law:

P1V1/T1 = P2V2/T2,

where P - pressure, V - volume, T - temperature.

As the pressure is constant:

V1/T1 = V2/T2.

From here:

V2 = V1T2/T1

Here, V1 = 2.50 L, T1 = -196°C = 77.15 K, T2 = 100°C = 373.15 K.

As a result,

V2 = 2.5 L × 373.15 K / 77.15 K = 12.1 L

Answer: 12.1 L.